# Supervised Graph Contrastive Learning for Gene Regulatory Networks

### v4

@@ F.6 Evaluation on Link Prediction
Appendix G Cross-Domain Training
Our proposed method ultimately aims to enable cancer-type–agnostic representation learning of GRNs. Since SupGCL is a scale-free graph representation learning framework with respect to the set of genes and the knockdown perturbations applied, it can in principle be applied to diverse GRNs derived from different cancer types. Accordingly, pre-training is feasible even when the patient graphs $\mathcal{G}$ and supervision graphs $\mathcal{H}_{a}$ come from different cancer types. Moreover, downstream tasks can be learned even if the cancer types used for pre-training and fine-tuning differ. It is also possible to take a model fine-tuned on one cancer type and directly apply it to predict outcomes for another. Nevertheless, the extent to which SupGCL retains predictive accuracy under such cross-domain conditions requires careful investigation.
To examine cancer-type dependency, we conducted three cross-domain experiments. (i) Pre-training phase: using knockdown GRNs from a different cancer type than the patient GRNs. (ii) Fine-tuning phase: applying a pre-trained model from one cancer type to fine-tune on another. (iii) OOD setting: applying a model pre-trained and fine-tuned on one cancer type directly to patients of a different type. The results are summarized in Table 18.
From experiment (i), we found that using breast knockdown data to supervise learning on lung or colorectal patient GRNs yielded performance comparable to the in-domain setting. In experiment (ii), performance was preserved when the model was pre-trained on lung data and fine-tuned on breast, but performance degraded when using colorectal as the source. A similar trend was observed in experiment (iii). Overall, these results suggest that SupGCL can generalize across cancer types at the pre-training stage, where the supervision GRN comes from a different cancer type. However, differences in cancer type during the fine-tuning phase cannot be ignored: models fine-tuned on a single cancer type do not exhibit robustness when directly applied to unseen cancer types.
@@ Appendix G Cross-Domain Training
cross domain 1: model Lung $\rightarrow$ Breast $0.654\pm 0.075$ $0.248\pm 0.043$
cross domain 2: model Colorectal $\rightarrow$ Breast $0.632\pm 0.072$ $0.249\pm 0.030$
@@
Appendix H Robustness Analysis of SupGCL
In this section, we evaluate the robustness of the proposed SupGCL method. We analyze three aspects: (a) scaling with respect to the number of pretraining samples, (b) the impact of the sample size of supervision graphs, and (c) robustness to noise in the estimated GRNs.
a. Scaling with Respect to Pretraining Data Size
@@ Appendix H Robustness Analysis of SupGCL
b. Effect of Supervision Graph Sample Size
@@
c. Robustness to Noise in Estimated GRNs
In our framework, edge contributions from estimated GRNs are used as input features to the GNN. To evaluate robustness against noise in estimated GRNs, we performed additional experiments.
We employed Bayesian network inference for GRN estimation (see Appendix D for details), which is known to be sensitive to initialization and may introduce errors. To mitigate this, we estimated GRNs 1000 times and applied frequency-based filtering of edges. The current setting applies a cutoff of 5%, but we additionally evaluated thresholds of 3% and 10%. The results (Table 21) show only minor differences across thresholds, confirming that SupGCL is robust to uncertainty inherent in GRN estimation.
Table 21: Performance under different filtering thresholds in GRN estimation (Breast Cancer GRN).
@@
SupGCL (3% threshold) $0.665\pm 0.059$ $0.238\pm 0.029$
SupGCL (5% threshold, original) $0.650\pm 0.059$ $0.243\pm 0.052$
SupGCL (10% threshold) $0.646\pm 0.039$ $0.244\pm 0.029$
Appendix I Biological Enrichment Analysis Using Latent Representations
@@
a. Clustering in the Latent Space
@@
b. Gene Ontology Enrichment Analysis
@@
c. KEGG Pathway Enrichment Analysis
c.1 KEGG Results
@@
  • Ashburner et al. (2000) Michael Ashburner, Catherine A Ball, Judith A Blake, David Botstein, Heather Butler, J Michael Cherry, Allan P Davis, Kara Dolinski, Selina S Dwight, Janan T Eppig, et al. Gene ontology: tool for the unification of biology. Nature genetics, 25(1):25–29, 2000.
  • Costa et al. (2010) Pedro R Costa, Marcio L Acencio, and Ney Lemke. A machine learning approach for genome-wide prediction of morbid and druggable human genes based on systems-level data. In BMC genomics, volume 11, pp. 1–15. Springer, 2010.
  • Friedman et al. (2000) Nir Friedman, Michal Linial, Iftach Nachman, and Dana Pe’er. Using Bayesian networks to analyze expression data. In Proceedings of the fourth annual international conference on Computational molecular biology, RECOMB ’00, pp. 127–135, New York, NY, USA, 2000. Association for Computing Machinery. ISBN 978-1-58113-186-4.
  • Imoto et al. (2003) Seiya Imoto, Sunyong Kim, Takao Goto, Sachiyo Aburatani, Kousuke Tashiro, Satoru Kuhara, and Satoru Miyano. Bayesian network and nonparametric heteroscedastic regression for nonlinear modeling of genetic network. Journal of bioinformatics and computational biology, 1(02):231–252, 2003.
  • Langfelder & Horvath (2008) Peter Langfelder and Steve Horvath. WGCNA: an R package for weighted correlation network analysis. BMC Bioinformatics, 9(1):559, December 2008. ISSN 1471-2105.
  • Margolin et al. (2006) Adam A. Margolin, Ilya Nemenman, Katia Basso, Chris Wiggins, Gustavo Stolovitzky, Riccardo Dalla Favera, and Andrea Califano. ARACNE: An Algorithm for the Reconstruction of Gene Regulatory Networks in a Mammalian Cellular Context. BMC Bioinformatics, 7(1):S7, March 2006. ISSN 1471-2105.
  • Paolacci et al. (2009) Anna R Paolacci, Oronzo A Tanzarella, Enrico Porceddu, and Mario Ciaffi. Identification and validation of reference genes for quantitative rt-pcr normalization in wheat. BMC molecular biology, 10:1–27, 2009.
  • Shu et al. (2021) Hantao Shu, Jingtian Zhou, Qiuyu Lian, Han Li, Dan Zhao, Jianyang Zeng, and Jianzhu Ma. Modeling gene regulatory networks using neural network architectures. Nature Computational Science, 1(7):491–501, July 2021. ISSN 2662-8457. Publisher: Nature Publishing Group.
@@
  • Tamada et al. (2011) Yoshinori Tamada, Teppei Shimamura, Rui Yamaguchi, Seiya Imoto, Masao Nagasaki, and Satoru Miyano. Sign: Large-Scale Gene Network Estimation Environment for High Performance Computing. Genome Informatics, 25(1):40–52, 2011.
  • Thakoor et al. (2021) Shantanu Thakoor, Corentin Tallec, Mohammad Gheshlaghi Azar, Mehdi Azabou, Eva L Dyer, Remi Munos, Petar Veličković, and Michal Valko. Large-scale representation learning on graphs via bootstrapping. arXiv preprint arXiv:2102.06514, 2021.
  • Tomoto et al. (2024) Mei Tomoto, Yohei Mineharu, Noriaki Sato, Yoshinori Tamada, Mari Nogami-Itoh, Masataka Kuroda, Jun Adachi, Yoshito Takeda, Kenji Mizuguchi, Atsushi Kumanogoh, Yayoi Natsume-Kitatani, and Yasushi Okuno. Idiopathic pulmonary fibrosis-specific Bayesian network integrating extracellular vesicle proteome and clinical information. Scientific Reports, 14(1):1315, January 2024. ISSN 2045-2322. Publisher: Nature Publishing Group.
@@
  • Xin et al. (2015) Jiwen Xin, Adam Mark, Cyrus Afrasiabi, Ginger Tsueng, Moritz Juchler, Nikhil Gopal, Gregory S Stupp, Timothy E Putman, Benjamin J Ainscough, Obi L Griffith, et al. Mygene. info and myvariant. info: gene and variant annotation query services. bioRxiv, pp. 035667, 2015.
  • Yahara et al. (2022) Hiroko Yahara, Souichi Yanamoto, Miho Takahashi, Yuji Hamada, Haruo Sakamoto, Takuya Asaka, Yoshimasa Kitagawa, Kuniyasu Moridera, Kazuma Noguchi, Masaya Sugiyama, Yutaka Maruoka, and Koji Yahara. Whole blood transcriptome profiling identifies gene expression subnetworks and a key gene characteristic of the rare type of osteomyelitis. Biochemistry and Biophysics Reports, 32:101328, December 2022. ISSN 2405-5808.
  • Zhu et al. (2021) Yanqiao Zhu, Yichen Xu, Feng Yu, Qiang Liu, Shu Wu, and Liang Wang. Graph contrastive learning with adaptive augmentation. In Proceedings of the Web Conference 2021, WWW ’21, pp. 2069–2080. ACM, April 2021.

### v5

@@ G.5 Evaluation on Link Prediction
Appendix H Cross-Domain Training
H.1 Cross-Domain Evaluation Results
Our proposed framework ultimately aims to enable cancer-type–agnostic representation learning of GRNs. SupGCL is scale-free with respect to both the gene set and the set of knockdown perturbations, and can in principle be applied to diverse GRNs derived from different cancer types. Accordingly, pre-training remains feasible even when the patient graphs $\mathcal{G}$ and the supervision graphs $\mathcal{H}_{a}$ originate from different cancer types. Downstream tasks can also be learned when the cancer types used for pre-training and fine-tuning differ, and it is further possible to apply a model fine-tuned on one cancer type directly to outcome prediction in another. However, the degree to which SupGCL maintains predictive accuracy under such cross-domain conditions requires careful empirical evaluation.
To investigate cancer-type dependency, we conducted three types of cross-domain experiments:
  1. (i)
    Pre-training phase: a setting in which knockdown GRNs from a different cancer type than the patient GRNs are used as supervision.
  2. (ii)
    Fine-tuning phase: a setting in which a model pre-trained on one cancer type is fine-tuned on data from another cancer type.
  3. (iii)
    OOD setting: a setting in which a model pre-trained and fine-tuned on one cancer type is directly applied to patient GRNs from an unseen cancer type.
The configurations and results for SupGCL are summarized in Table 20.
In addition, to examine whether this behavior is specific to SupGCL or common to GCL methods more broadly, we performed the same cross-domain analysis with GRACE, a representative general-purpose GCL model. For a fair comparison, we aligned the settings for (ii) cross-domain fine-tuning and (iii) OOD application with those used for SupGCL and evaluated both methods under identical conditions. The corresponding results for GRACE are reported in Table 21.
For SupGCL, experiment (i) shows that using breast knockdown data as supervision for lung or colorectal patient GRNs yields performance comparable to the in-domain setting, indicating that SupGCL can generalize across cancer types at the pre-training stage when only the teacher GRN domain is changed. In experiment (ii), performance is largely preserved when using lung as the pre-training source and breast as the target for fine-tuning, whereas using colorectal as the source leads to a noticeable degradation. A similar trend is observed in experiment (iii), where directly applying a breast-trained model to lung or colorectal patients results in a substantial drop in performance.
[…]
Taken together, these results indicate that SupGCL can maintain reasonable generalization when using knockdown GRNs from different cancer types during pre-training, but—much like GRACE—suffers from performance deterioration when the cancer type changes at the fine-tuning and inference stages. In particular, models fine-tuned on a single cancer type lack robustness when directly applied to unseen cancer types, highlighting the need for learning strategies that more explicitly promote cross-cancer generalization.
@@ H.1 Cross-Domain Evaluation Results
[…]
@@
Table 21: Cross-domain performance of GRACE
Setting Configuration Hazard (c-index) BP (Subset Acc.)
[…]
(iii) OOD: cancer type of fine-tuning / prediction
original: Lung / Lung $0.609\pm 0.055$ $0.259\pm 0.063$
cross domain 1: Breast / Lung $0.534$ $0.148$
original: Colorectal / Colorectal $0.647\pm 0.059$ $0.249\pm 0.050$
cross domain 2: Breast / Colorectal $0.562$ $0.133$
H.2 Toward Pan-Cancer Representation Learning
The results in Section H.1 show that both SupGCL and GRACE suffer from performance degradation when the cancer type differs at the fine-tuning and inference stages. This suggests that pre-training on a single cancer type has inherent limitations in terms of generalizability and motivates learning strategies that incorporate more diverse biological backgrounds.
To explore the potential of more cancer-type–agnostic representations, we conducted a preliminary pan-cancer pre-training experiment in which GRNs from three cancer types (breast, lung, and colorectal) were jointly used during pre-training. In this setting, only the pre-training phase is multi-cancer (pan-cancer), while fine-tuning and downstream evaluation are restricted to breast cancer data. We then compared this pan-cancer pre-training with the single-cancer pre-training model that uses only breast GRNs.
The results are summarized in Table 22. We found that pan-cancer pre-training does not cause a substantial drop in performance compared with breast-only pre-training. This indicates that integrating multiple cancer types at the pre-training stage can retain, and potentially enrich, the learned representations without harming downstream performance on a specific cancer type. We expect that extending this approach to include a broader set of cancer types and combining it with domain adaptation techniques could further enhance cross-cancer generalization in future work.
Table 22: Pan-cancer pre-training vs. single-cancer pre-training (downstream task: breast cancer).
Model Setting Hazard (c-index) BP (Subset Accuracy)
Breast $\rightarrow$ Breast $0.650\pm 0.059$ $0.243\pm 0.052$
Pan-cancer $\rightarrow$ Breast $0.649\pm 0.048$ $0.241\pm 0.052$
Appendix I Robustness Analysis of SupGCL
In this section, we evaluate the robustness of the proposed SupGCL method. We analyze three aspects: 1. scaling with respect to the number of pretraining samples, 2. the impact of the sample size of supervision graphs, and 3. robustness to noise in the estimated GRNs.
I.1 Scaling with Respect to Pretraining Data Size
@@
I.2 Effect of Supervision Graph Sample Size
@@
I.3 Robustness to Noise in Estimated GRNs
In our framework, edge contribution scores derived from estimated GRNs are used as edge features for the GNN. To evaluate how robust the proposed method is to noise in these estimated GRNs, we conducted additional experiments.
We adopt Bayesian network inference for GRN estimation (see Appendix D for details), which is known to be sensitive to initialization and can introduce errors. To mitigate this, we estimate GRNs 1000 times and apply frequency-based filtering of edges to reduce noise. In the main setting, we apply a uniform cutoff of 5% for both the patient GRN $\mathcal{G}$ and the teacher GRN $\mathcal{H}$. For comparison, we additionally evaluate thresholds of 3% and 10% for both graphs, and also consider an asymmetric configuration in which we retain only higher-confidence edges in the patient GRN while allowing lower-confidence edges to remain in the teacher GRN, i.e., $(\mathcal{G}:10\%,\mathcal{H}:3\%)$. The results are summarized in Table 25, and the statistics of each GRN are presented in Table 26.
Overall, differences in downstream performance across cutoff values are small, indicating that SupGCL can pretrain robustly despite uncertainty inherent in GRN estimation. Moreover, even when the teacher GRN contains a non-negligible amount of noise (e.g., in the $(\mathcal{G}:10\%,\mathcal{H}:3\%)$ setting), the performance of SupGCL does not substantially degrade, suggesting that the method also exhibits robustness to noise in the perturbation graphs used as supervision.
Table 25: Downstream performance under different filtering thresholds in GRN estimation (breast cancer GRN).
@@ I.3 Robustness to Noise in Estimated GRNs
SupGCL ($\mathcal{G}$:3%, $\mathcal{H}$:3%) $0.665\pm 0.059$ $0.238\pm 0.029$
SupGCL ($\mathcal{G}$:5%, $\mathcal{H}$:5%, original) $0.650\pm 0.059$ $0.243\pm 0.052$
SupGCL ($\mathcal{G}$:10%, $\mathcal{H}$:10%) $0.646\pm 0.039$ $0.244\pm 0.029$
SupGCL ($\mathcal{G}$:10%, $\mathcal{H}$:3%) $0.649\pm 0.055$ $0.242\pm 0.045$
Table 26: Statistics of GRNs constructed with different edge cutoff thresholds
3% 5% 10%
TCGA LINCS TCGA LINCS TCGA LINCS
number of nodes 975 975 975 975 975 975
number of edges 15405 12057 13170 11209 10201 10061
average degree 15.8 12.366154 13.507692 11.49641 10.462564 10.318974
Appendix J Biological Enrichment Analysis Using Latent Representations
@@
J.1 Clustering in the Latent Space
@@
J.2 Gene Ontology Enrichment Analysis
@@
J.3 KEGG Pathway Enrichment Analysis
J.3.1 KEGG Results
@@
Appendix K Comparison with Other Models
Table 30: Comparison with additional models (breast cancer). Geneformer condition # 1: freeze all except last layer, # 2: freeze first 2 layers, # 3: full fine-tuning.
Model Hazard (C-index) BP (Subset Accuracy)
GCA $0.620\pm 0.039$ $0.241\pm 0.021$
AFGRL $0.616\pm 0.037$ $0.240\pm 0.026$
SimGRACE $0.638\pm 0.032$ $0.228\pm 0.044$
AD-GCL $0.636\pm 0.036$ $0.228\pm 0.023$
AutoGCL $0.620\pm 0.058$ $0.214\pm 0.021$
ArieL $0.626\pm 0.041$ $0.230\pm 0.039$
Geneformer # 1 $0.604\pm 0.096$ $0.241\pm 0.076$
Geneformer # 2 $0.595\pm 0.054$ $0.230\pm 0.057$
Geneformer # 3 $0.585\pm 0.048$ $0.226\pm 0.066$
Muse-GNN (adapted) $0.639\pm 0.046$ $0.238\pm 0.037$
SupGCL (Ours) $\mathbf{0.650\pm 0.059}$ $\mathbf{0.243\pm 0.052}$
K.1 Comparison with other graph contrastive learning models
In graph contrastive learning (GCL), topological changes introduced by augmentations such as node or edge dropping are known to degrade performance. To address this issue, recent works have proposed frameworks that either correct or avoid the effects of augmentation. For instance, SGRL (He et al., 2024), the successor to BGRL (Thakoor et al., 2021), has established a state-of-the-art augmentation-free paradigm. Other representative methods include GCA (Zhu et al., 2021), which automatically adjusts and corrects augmentation effects; AFGRL (Lee et al., 2022), which integrates reconstruction and bootstrap learning; and SimGRACE (Xia et al., 2022a), which perturbs model parameters with simple Gaussian noise.
Beyond these, more advanced frameworks explicitly treat augmentations as learnable objects. AD-GCL (Suresh et al., 2021) adopts a min–max adversarial optimization scheme in which the encoder and a perturbation generator compete to search for destructive transformations. AutoGCL (Yin et al., 2022) introduces a learnable policy generator that selects node-wise drop/mask/keep operations, and ArieL (Feng et al., 2024) employs Projected Gradient Descent (PGD) to construct adversarial views together with an information regularization term. All of these methods share the common philosophy of not fixing the augmentation design in advance, but instead optimizing augmentations jointly with the encoder. However, the perturbations learned in these frameworks are not semantic operations grounded in real-world mechanisms; they remain generic regularization noise, so the amount of data that explains or constrains the augmentations themselves is not increased.
In contrast, SupGCL differs fundamentally from augmentation-free and augmentation-adjustment approaches. Rather than avoiding topological changes, SupGCL leverages them as informative signals by supervising augmentations with knockdown-derived GRNs. That is, while conventional approaches aim to circumvent or correct augmentation-induced topology shifts, SupGCL explicitly exploits them as positive supervision, transforming a traditionally negative factor into a beneficial learning signal.
Table 30 reports the comparison on breast cancer GRNs for hazard prediction (graph-level) and BP classification (node-level). As shown in the results, SupGCL consistently outperforms representative augmentation-correction and augmentation-free models (GCA, AFGRL, SimGRACE, AD-GCL, AutoGCL, ArieL). This suggests that actively learning from topology changes as supervised signals, grounded in knockdown GRNs, leads to improved representation learning performance.
K.2 Comparison with biological foundation model (Geneformer)
Recent years have seen rapid progress in the development of biological foundation models. A representative example is Geneformer (Theodoris et al., 2023), a transformer-based foundation model pretrained on tens of millions of single-cell transcriptomes. Although the data modality and task setup differ from those in this work, both SupGCL and Geneformer ultimately aim to learn useful representations of genes. Therefore, comparing SupGCL against such a foundation model is important for positioning and validating our approach.
Geneformer converts each cell’s gene expression vector into a sequence of gene tokens ordered by expression rank, and then uses a Transformer to learn context-dependent representations of genes and cells. In this study, we use the publicly available default model Geneformer-316M (embedding dimension 1152, 18 transformer layers) and finetune it on our task settings. Specifically, we consider three finetuning regimes:
  1. (i)
    updating only the final transformer layer and the prediction head,
  2. (ii)
    partially finetuning by freezing only the first two encoder layers while updating the remaining layers and the prediction head
  3. (iii)
    full finetuning, where all layers are updated.
For the node-level task, we feed each patient’s bulk expression profile into Geneformer and use the resulting gene embeddings for multi-label GO-BP classification. For the graph-level tasks, we construct patient representations by pooling the CLS token and apply them to breast cancer subtype classification and hazard prediction.
Table 30 summarizes the comparison between SupGCL and Geneformer. On the node-level gene classification task, Geneformer achieves reasonable performance when updating only the last layer or under partial finetuning, confirming the representational power of a transformer-based foundation model. However, for the graph-level tasks of patient prognosis prediction and subtype classification, SupGCL consistently outperforms Geneformer under all finetuning settings, with particularly poor performance observed under full finetuning. This degradation is likely attributable to a distributional mismatch, as Geneformer is designed and pretrained for single-cell transcriptomes, as well as to its inability to directly exploit patient-specific GRN structure.
Overall, while Geneformer—a large-scale transformer foundation model trained on single-cell data—provides competitive performance for node-level functional label prediction, it is outperformed by SupGCL on the GRN-based graph-level tasks that are the focus of this work. In particular, SupGCL’s direct use of knockdown GRNs as supervision appears crucial for achieving strong performance on patient-level outcomes.
K.3 Comparison with MuSeGNN (adapted)
We also compared SupGCL with MuSe-GNN (Liu et al., 2023), a representative multi-omics graph framework. While MuSe-GNN typically integrates distinct omics modalities (e.g., scRNA-seq, scATAC-seq, spatial transcriptomics), we adapted it to our setting. Note that this experimental setup utilizes data types that differ from the specific multi-omics modalities originally assumed by MuSe-GNN, but maintains the core principle of multi-view representation learning.
Specifically, we employed the patient GRNs $\mathcal{G}$ derived from TCGA bulk RNA-seq and the teacher knockdown GRNs $\mathcal{H}^{*}$ derived from LINCS experiments as the dual input sources. To rigorously implement this comparison, we constructed a MuSe-GNN-adapted architecture with independent encoders $f_{\mathcal{G}}$ and $f_{\mathcal{H}}$ for each domain:
$f_{\mathcal{G}}:\mathcal{G}_{a}\mapsto Z^{a}=\{z_{i}^{a}\}_{i},\quad f_{\mathcal{H}}:\mathcal{H}_{a}\mapsto Y^{a}=\{y_{i}^{a}\}_{i}$
In the original MuSe-GNN framework, the total loss is defined as a combination of three components:
$\mathrm{Loss}_{\rm MuSe-GNN}\triangleq\mathrm{Loss}_{\rm Recon}+\mathrm{Loss}_{\rm Common\text{-}Sim}+\mathrm{Loss}_{\rm Diff\text{-}Sim},$ (43)
which correspond to (i) edge reconstruction loss, (ii) similarity learning for genes common to both modalities, and (iii) contrastive learning for genes specific to one modality.
In our experimental setting, the node sets for $\mathcal{G}$ and $\mathcal{H}^{*}$ are identical. Consequently, the modality-specific term $\mathrm{Loss}_{\rm Diff\text{-}Sim}$ is not applicable. We therefore optimized the model using $\mathrm{Loss}_{\rm Recon}$ and $\mathrm{Loss}_{\rm Common\text{-}Sim}$. For the reconstruction term, we applied the standard graph reconstruction loss to both domains. For the similarity term, we adopted a standard InfoNCE formulation to promote the alignment of the corresponding node embeddings $z_{i}^{a}$ and $y_{i}^{a}$ across the two domains.
The results are summarized in Table 30. Although the adapted MuSe-GNN demonstrated high performance in both graph-level and node-level tasks, it did not surpass SupGCL. These results suggest that our proposed setting—where the knockdown GRN $\mathcal{H}$ provides a teacher distribution for the learner GRN $\mathcal{G}$—allows SupGCL to formulate the structural relationship more directly and validly than treating them as multi-modal inputs.
Appendix L Augmentation in GCL
L.1 Meaning of Augmentation Learning of GCL
In this study, we adopted a supervised approach where biological priors determine the graph perturbations. A natural counter-argument might suggest optimizing the graph augmentation strategy itself—learning to generate the ”optimal” augmented graph rather than using fixed rules. Here, we theoretically analyze why learning the augmentation model within our contrastive framework leads to trivial solutions, justifying our choice of fixed biological supervision.
Our proposed SupGCL framework is rigorously grounded in the minimization of the KL divergence between the joint distributions of node pairs $(i,j)$ and augmentation pairs $(a,b)$. The loss function is defined as:
$\mathrm{Loss}_{\rm GCL}=D_{\mathrm{KL}}(p_{\phi}(i,j,a,b)\mid q_{\phi}(i,j,a,b)).$
Hereafter, we define $a,b\in\mathcal{K}$ as the indices of the target genes subjected to knockdown experiments, rather than abstract augmentation indices. This allows us to formulate the augmentation generation process as follows.
To discuss “optimal graph augmentation” in this context, one must consider modeling a virtual graph augmentation generator $q_{\psi}(\mathcal{G}_{a}\mid\mathcal{G},a)$, parameterized by $\psi$. Here, $\mathcal{G}_{a}$ denotes the generated graph augmentation corresponding to the knockdown index $a$, which aims to approximate the biological ground truth $\mathcal{H}_{a}$. The objective would be to learn this generator $q_{\psi}$ while simultaneously minimizing the KL divergence over the knockdown set $\mathcal{K}$ and node set $\mathcal{V}$.
However, if the graph augmentation model $q_{\psi}(\mathcal{G}_{a}\mid\mathcal{G},a)$ is not fixed and is trained jointly with the representation parameters $\phi$, the optimization becomes ill-posed. Specifically, the reference distribution $p_{\phi}(i,j,a,b)$ and the target distribution $q_{\psi,\phi}(i,j,a,b)$ would be optimized via separate parameters without sufficient constraints. This formulation allows the model to converge to a trivial solution.
A representative example of such a trivial solution is the mapping $g_{\psi}(\mathcal{G},a)=\mathcal{H}_{a},~\forall\mathcal{G}$. In this scenario, the generator learns to output the teacher graph $\mathcal{H}_{a}$ directly, completely ignoring the structural information of the input patient graph $\mathcal{G}$. While this minimizes the distributional distance, it results in a ”data extension” that is independent of the input graph $\mathcal{G}$. Consequently, the learned representations would fail to capture the unique topological features of the patient-specific networks, rendering them useless for downstream tasks that require patient-specific biological insights.
Therefore, we argue that a mathematical formulation seeking ”contrastively effective data augmentation” often results in a ”contrastive formulation capable of trivial solutions.” Such solutions may be mathematically valid in an unconstrained optimization landscape but are unsuitable for downstream tasks requiring biological fidelity.
While it is technically possible to prevent this collapse by introducing additional constraints—such as reconstruction losses via AutoEncoders or regularization on the weights of $q_{\psi}$—our goal in this work was to establish the simplest and most interpretable formulation of supervised contrastive learning. Thus, we deliberately excluded learned augmentations in favor of a probabilistic formulation that directly leverages biological experimental data as explicit supervision.
L.2 Comparison of Data Augmentation Strategies
In this work, we adopt an artificial augmentation scheme in which the node features of a perturbed gene are set to zero and all its incident edges are deleted. The motivation is to mimic the effect of a gene knockdown, which reduces the expression level of a specific gene and weakens its interactions with other genes. Indeed, prior studies on virtual knockdown and in silico gene perturbation employ similar strategies: CellOracle (Yang et al., 2023) simulates genetic perturbations by replacing the expression of the target gene with zero, while GenKI (Kamimoto et al., 2023) and scTenifoldKnk (Osorio et al., 2022) model knockdown effects on GRNs by removing all edges adjacent to the perturbed gene. Our artificial augmentation design follows these precedents.
In addition to this default augmentation, we further evaluated several alternative perturbation modes on the breast cancer dataset: (i) node copy: replacing the node features of the target gene with those of another randomly chosen gene; (ii) in-edge: deleting only the incoming edges of the target gene; and (iii) out-edge: deleting only its outgoing edges. The results are summarized in Table 31.
Overall, the node-copy scheme (i) yields slightly worse performance than the others, but the differences across perturbation modes are modest. One plausible interpretation is that, whereas knockdown corresponds to a decrease in expression, randomly copying features from another gene can sometimes increase the effective expression of the target gene, thereby inducing perturbations that are misaligned with the underlying biological process. These observations suggest that augmentations that more faithfully reflect the biology of knockdown-like perturbations tend to support better representation learning.
Table 31: Performance comparison under different perturbation modes (breast cancer GRN).
Method Hazard (c-index) BP (Subset Accuracy)
node copy $0.647\pm 0.052$ $0.236\pm 0.026$
in-edge $0.653\pm 0.077$ $0.242\pm 0.039$
out-edge $0.667\pm 0.059$ $0.240\pm 0.042$
original setting $0.650\pm 0.059$ $0.243\pm 0.052$
@@
  • Ashburner et al. (2000) Ashburner, M., Ball, C. A., Blake, J. A., Botstein, D., Butler, H., Cherry, J. M., Davis, A. P., Dolinski, K., Dwight, S. S., Eppig, J. T., et al. Gene ontology: tool for the unification of biology. Nature genetics, 25(1):25–29, 2000.
  • Costa et al. (2010) Costa, P. R., Acencio, M. L., and Lemke, N. A machine learning approach for genome-wide prediction of morbid and druggable human genes based on systems-level data. In BMC genomics, volume 11, pp. 1–15. Springer, 2010.
  • Friedman et al. (2000) Friedman, N., Linial, M., Nachman, I., and Pe’er, D. Using Bayesian networks to analyze expression data. In Proceedings of the fourth annual international conference on Computational molecular biology, RECOMB ’00, pp. 127–135, New York, NY, USA, 2000. Association for Computing Machinery. ISBN 978-1-58113-186-4.
  • Imoto et al. (2003) Imoto, S., Kim, S., Goto, T., Aburatani, S., Tashiro, K., Kuhara, S., and Miyano, S. Bayesian network and nonparametric heteroscedastic regression for nonlinear modeling of genetic network. Journal of bioinformatics and computational biology, 1(02):231–252, 2003.
  • Langfelder & Horvath (2008) Langfelder, P. and Horvath, S. WGCNA: an R package for weighted correlation network analysis. BMC Bioinformatics, 9(1):559, December 2008. ISSN 1471-2105.
  • Margolin et al. (2006) Margolin, A. A., Nemenman, I., Basso, K., Wiggins, C., Stolovitzky, G., Favera, R. D., and Califano, A. ARACNE: An Algorithm for the Reconstruction of Gene Regulatory Networks in a Mammalian Cellular Context. BMC Bioinformatics, 7(1):S7, March 2006. ISSN 1471-2105.
  • Paolacci et al. (2009) Paolacci, A. R., Tanzarella, O. A., Porceddu, E., and Ciaffi, M. Identification and validation of reference genes for quantitative rt-pcr normalization in wheat. BMC molecular biology, 10:1–27, 2009.
  • Shu et al. (2021) Shu, H., Zhou, J., Lian, Q., Li, H., Zhao, D., Zeng, J., and Ma, J. Modeling gene regulatory networks using neural network architectures. Nature Computational Science, 1(7):491–501, July 2021. ISSN 2662-8457. Publisher: Nature Publishing Group.
@@
  • Tamada et al. (2011) Tamada, Y., Shimamura, T., Yamaguchi, R., Imoto, S., Nagasaki, M., and Miyano, S. Sign: Large-Scale Gene Network Estimation Environment for High Performance Computing. Genome Informatics, 25(1):40–52, 2011.
  • Thakoor et al. (2021) Thakoor, S., Tallec, C., Azar, M. G., Azabou, M., Dyer, E. L., Munos, R., Veličković, P., and Valko, M. Large-scale representation learning on graphs via bootstrapping. arXiv preprint arXiv:2102.06514, 2021.
  • Theodoris et al. (2023) Theodoris, C. V., Xiao, L., Chopra, A., Chaffin, M. D., Al Sayed, Z. R., Hill, M. C., Mantineo, H., Brydon, E. M., Zeng, Z., Liu, X. S., et al. Transfer learning enables predictions in network biology. Nature, 618(7965):616–624, 2023.
  • Tomoto et al. (2024) Tomoto, M., Mineharu, Y., Sato, N., Tamada, Y., Nogami-Itoh, M., Kuroda, M., Adachi, J., Takeda, Y., Mizuguchi, K., Kumanogoh, A., Natsume-Kitatani, Y., and Okuno, Y. Idiopathic pulmonary fibrosis-specific Bayesian network integrating extracellular vesicle proteome and clinical information. Scientific Reports, 14(1):1315, January 2024. ISSN 2045-2322. Publisher: Nature Publishing Group.
@@
  • Xin et al. (2015) Xin, J., Mark, A., Afrasiabi, C., Tsueng, G., Juchler, M., Gopal, N., Stupp, G. S., Putman, T. E., Ainscough, B. J., Griffith, O. L., et al. Mygene. info and myvariant. info: gene and variant annotation query services. bioRxiv, pp. 035667, 2015.
  • Yahara et al. (2022) Yahara, H., Yanamoto, S., Takahashi, M., Hamada, Y., Sakamoto, H., Asaka, T., Kitagawa, Y., Moridera, K., Noguchi, K., Sugiyama, M., Maruoka, Y., and Yahara, K. Whole blood transcriptome profiling identifies gene expression subnetworks and a key gene characteristic of the rare type of osteomyelitis. Biochemistry and Biophysics Reports, 32:101328, December 2022. ISSN 2405-5808.
  • Zhu et al. (2021) Zhu, Y., Xu, Y., Yu, F., Liu, Q., Wu, S., and Wang, L. Graph contrastive learning with adaptive augmentation. In Proceedings of the Web Conference 2021, WWW ’21, pp. 2069–2080. ACM, April 2021.